Monday 7th November 2016

  • Adjustments for SEDM on campus during March-April, coinciding with the end of iPTF.
  • Shipping container for transporting SEDM on unknown location. Contacted Nick to figure out where the container is located.
  • The upgrade shall focus on the following points:
    • Stabilize the optics. There are some springs that do not look very stretched out. There is also a lens that is re-coated in the right side. Changing that lens would imply to re-align the optics. The optical components are in compact boxes, so there shall be a strong reason to change this lens, as the alignment shall take a lot of time. Something to consider if there is extra time left. Option to bring Nick as a consultant for this operation.
    • Move electronics outside of the box to stabilize the temperature. Also add some additional baffling. 3D printing facilities at Caltech could be used for that. Moving electronics outside of the box shall allow to organize all cables better.
    • Seal the enclosure. Most enclosure holes are unusued, so sealing them would provide better sealing again dust. Nitrogen seal could be possible, given that the P48 has it. A daily flow shall be enough and have low maintenance. Currently there is a humidity absorbent, but it is completely saturated because it is not well isolated. The enclosure could be redesigned, so that cables should be moved outside. There seems to be a lot of extra space in the instrument. Creating a new enclosure should not take more than 1-2 weeks, so that it is viable to have a new one by March-April. Mechanical safety shall be considered, as we should not leave any cameras outside of a protective box.
    • Improve the focus. Currently there are some weak springs that are not design to push against the focus nudge. Better springs are available. The return force is not stable right now, so that gravity is one of the factors that can affect the current focus.
  • New CAD design will be created to reflect all this new improvements. A total amount of 6 weeks downtime shall be good to accommodate all this changes.
  • Baffling for the RC filters seems to be non-trivial, so that a rather more complex data solution would be needed.

We need a thermal controller or some sensors are about to be ordered, so that the thermal sensor that could be put inside in ~1 month time. No need to wait until the general adjustments on campus.

Pointing position

  • The error in position seems to have improved, but it is yet not perfect. The telescope was not X-ed properly though.
  • We need a better analysis on the offsets. Colour code by date shall be a good indication.

-- NadiaBlagorodnova - 07 Nov 2016

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions
Topic revision: r2 - 2016-11-09 - RestonNash
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback